OpenStudy (anonymous):

olve 2x2 + 3x + 5 = 0. Round solutions to the nearest hundredth. x ≈ −2.14 and x ≈ 0.64 x ≈ −0.64 and x ≈ 2.14 x ≈ −4.28 and x ≈ 1.28 No real solutions i got c am i correct?

OpenStudy (ranga):

Discriminant is b^2 - 4ac (This is the part that is inside the square root when you use the formula for finding roots to a quadratic equation). If discriminant is negative it means you have to take the square root of a negative number which implies there are no real solutions. Is the discriminant positive or negative here?

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

or negative?

OpenStudy (anonymous):

-31

OpenStudy (ranga):

If it is negative then no real solution exists for this equation.

OpenStudy (anonymous):

so this is no solutions?

OpenStudy (ranga):

No REAL solutions. (But complex or imaginary solution does exist for this problem. But no REAL solution.)

OpenStudy (anonymous):

ok thank youu

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Can you help me with another please?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

I can try.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Identify the graph of a quadratic equation with two identical rational solutions.

OpenStudy (anonymous):

The graph of which quadratic equation is shown below?

OpenStudy (anonymous):

and that one

OpenStudy (ranga):

The roots or solutions of a quadratic equation on a graph is where the curve intersects the x axis: If the curve cuts the x axis at two points it means two different solutions. If the curve cuts the x axis at just one point it means two identical solutions. If the curve does not cut the x axis at all it means no real solutions So which one do you think is the answer?

OpenStudy (anonymous):

its either c or d

OpenStudy (anonymous):

d?

OpenStudy (ranga):

No. They are asking for the graph that shows two IDENTICAL rational solutions. It means the graph has to cut (or touch actually) at just ONE point.

OpenStudy (anonymous):

OHHH its B

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

What about the second one

OpenStudy (ranga):

What are the answer choices?

OpenStudy (anonymous):

y = x^2 + 2x − 8 y = x^2 − 2x − 8 y = −x^2 − 2x + 8 y = −x^2 − 4x − 8

OpenStudy (ranga):

The general quadratic equation is: ax^2 + bx + c = 0 If a is positive, the graph will open upward. If a is negative, the graph will open downward. So right away you can eliminate two possibilities. Then notice that when x = 0, y = +8 in the graph. Which of the remaining two choices satisfy that condition?

OpenStudy (anonymous):

the third one

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you so much for your help and Happy Halloween

OpenStudy (ranga):

you are welcome. Happy Halloween to you too.